# Orient Anything

### v2

@@ 3.1 Orientation estimation under rotational symmetries
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$. Let $\Omega_{S}$ be the shape’s orientation, and suppose $f^{*}:\mathcal{S}\rightarrow SO(3)$ is a solution to the following quotient regression problem:
@@
Then for any $R\in SO(3)$, $f^{*}(RS)=RQ^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
@@ 3.4 Implementation
We also employ test-time augmentation to improve our flipper model’s predictions. Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
@@ 4.2 Full-orientation estimation
We benchmark our pipeline’s performance (“Orient Anything”) against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation. Orient Anything achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis. In practice, one does not have access to this front-axis at inference time, so this represents a loose upper bound on the potential performance of a shape orientation algorithm built upon Upright-Net’s predicted up-axes.
@@
Orient Anything 0.00856 $\pm$ 0.0396
Orient Anything-APS 0.00208 $\pm$ 0.01407
@@
By applying each flip to the first-stage orienter’s output, we obtain a shape orientation pipeline that outputs sets of candidate shapes $\hat{S}_{c}$ given a single arbitrarily-rotated input $RS$. We call this method “Orient Anything-APS.” To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances. As one may trivially reduce the minimum chamfer distance over an APS by outputting arbitrarily large sets, we report a histogram of APS sizes in Figure 8 to demonstrate that these sets are typically small.
We report the performance of “Orient Anything-APS” in the final row of Table 2. Allowing our pipeline to output an APS reduces the mean chamfer distance between the closest shape in this set to the ground truth by a factor of 4 relative to Orient Anything without this feature. Figure 8 shows that with our chosen coverage probability, most APSs are small; the median APS size is 2, and 90% of APSs have at most 8 shapes. This demonstrates that our flipper typically assigns most of its probability mass to a few octahedral symmetries, which often contain a rotation that returns the shape to its canonical pose in Shapenet. In Figure 10 in Appendix C, we also include a grid of meshes in the prediction sets associated with 20 randomly-selected validation shapes. These prediction sets typically consist of a small collection of plausible shape orientations; by manually inspecting them, a human in the loop may select the orientation that is most suitable for their target application.
@@ 5 Conclusion
Impact statement.
@@
Acknowledgments.
We thank the authors of Pang et al. (2022) for their gracious assistance with their source code and Justin Solomon for valuable discussions. CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
@@ B.1 Quotient orienter
For test-time augmentation, we (1) randomly rotate the inputs $RS$ by $K$ random rotations $R_{k}\sim U(SO(3))$, $k=1,...,K$, (2) obtain the quotient orienter’s predictions $f_{\theta}(R_{k}RS)$ for each shape, (3) return these predictions to the original input’s orientation by computing $R_{k}^{\top}f_{\theta}(R_{k}RS)$, and (4) output the prediction $R_{k^{*}}^{\top}f_{\theta}(R_{k^{*}}RS)$ with the smallest average quotient distance to the remaining predictions.
@@ B.2 Flipper
We also employ test-time augmentation to improve our flipper model’s predictions. Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.

### v3

# Symmetry-Robust 3D Orientation Estimation

@@ 3.1 Orientation estimation under rotational symmetries
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$. Let $\Omega_{S}$ be the shape’s orientation, and consider the following quotient regression problem:
@@
Then for any $R\in SO(3)$, there exists a solution $f^{*}:\mathcal{S}\rightarrow SO(3)$ of the form $f^{*}(RS)=RQ^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
@@ 3.4 Implementation
We also employ test-time augmentation (TTA) to improve our flipper model’s predictions. Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
@@ 4.2 Full-orientation estimation
We benchmark our pipeline’s performance against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation. Our method achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis. In practice, one does not have access to this front-axis at inference time, so this represents a loose upper bound on the potential performance of a shape orientation algorithm built upon Upright-Net’s predicted up-axes.
@@
Ours (TTA) 0.00856 $\pm$ 0.0396
Ours (w/o TTA) 0.01107 $\pm$ 0.04342
Ours (with APS) 0.00208 $\pm$ 0.01407
@@
By applying each flip to the first-stage orienter’s output, we obtain a shape orientation pipeline that outputs sets of candidate shapes $\hat{S}_{c}$ given a single arbitrarily-rotated input $RS$. To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances. As one may trivially reduce the minimum chamfer distance over an APS by outputting arbitrarily large sets, we report a histogram of APS sizes in Figure 8 to demonstrate that these sets are typically small.
We report our pipeline’s performance with adaptive prediction sets in the last row of Table 2. Allowing our pipeline to output an APS reduces the mean chamfer distance between the closest shape in this set to the ground truth by a factor of 4 relative to Orient Anything without this feature. Figure 8 shows that with our chosen coverage probability, most APSs are small; the median APS size is 2, and 90% of APSs have at most 8 shapes. This demonstrates that our flipper typically assigns most of its probability mass to a few octahedral symmetries, which often contain a rotation that returns the shape to its canonical pose in Shapenet. In Figure 11 in Appendix C, we also include a grid of meshes in the prediction sets associated with 20 randomly-selected validation shapes. These prediction sets typically consist of a small collection of plausible shape orientations; by manually inspecting them, a human in the loop may select the orientation that is most suitable for their target application.
@@ 5 Conclusion
Acknowledgments
We thank the authors of Pang et al. (2022) for their gracious assistance with their source code and Justin Solomon for valuable discussions. Chris thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship, and gratefully acknowledges the support of a 2024 Exponent fellowship.
Impact statement
@@ B.1 Quotient orienter
Test-time augmentation for the quotient orienter.
Our use of TTA for the quotient orienter is motivated by the observation that the quotient orienter succeeds on most rotations and fails on small subsets of rotations. Given an arbitrarily-oriented input shape $RS$, we would like to mitigate the possibility that the shape is in an orientation for which the quotient orienter fails. To do so, we apply $K$ random rotations $R_{k}$ to the input shape to obtain randomly re-rotated shapes $R_{k}RS$. Because the quotient orienter only fails on small subsets of rotations, we expect it to succeed on most of the $R_{k}RS$ and recover the orientations of $R_{k}RS$ up to an octahedral symmetry.
Because we are interested in the orientation of $RS$ rather than $R_{k}RS$, we then need to apply the inverse of $R_{k}$ to each predicted orientation $f_{\theta}(R_{k}RS)$ to obtain a set of $K$ predicted orientations $R_{k}^{\top}f_{\theta}(R_{k}RS)$. These will be correct orientations of $RS$ up to octahedral symmetries for each $k$ where the quotient orienter succeeds.
Because we expect this to be the case for most $k$, we employ a voting scheme to select one of the candidate orientations $R_{k}^{\top}f_{\theta}(R_{k}RS)$. In this step, we compute each candidate’s average quotient $L_{2}$ loss with respect to every other candidate (i.e. the loss in Problem 3), and choose the candidate which minimizes this measure. Because we expect the quotient orienter to have succeeded for most of the $R_{k}RS$, this average quotient loss will be small for most candidates and large for the few outlier candidates on which the orienter failed. Choosing the candidate with the minimum average quotient loss with respect to the other candidates filters out these outliers and makes it likelier that we output the correct orientation of $RS$ up to an octahedral symmetry.
@@ B.2 Flipper
Test-time augmentation for the flipper.
The TTA procedure for the flipper is similar to the quotient orienter’s TTA procedure. We are given an input shape $FS$, which is correctly oriented up to an octahedral symmetry (a “flip”) $F$ if the quotient orienter succeeded. We apply $K$ random flips $F_{k}$ to the input shape to obtain randomly re-rotated shapes $F_{k}FS$ and apply the flipper to each of these shapes. Similarly to the quotient orienter, if the flipper succeeds on some $F_{k}FS$, it predicts $F_{k}F$ rather than the flip $F$ we are actually interested in. We consequently left-multiply each prediction $g(F_{k}FS)$ by $F_{k}^{\top}$, which maps each successful prediction $g(F_{k}FS)=F_{k}F$ to the true flip $F$. Because we expect the flipper to have succeeded on most inputs $F_{k}FS$ and failed on a minority of inputs, we again use a voting scheme to pick out the plurality prediction. In this case, we simply return the most common flip among the set of $F_{k}^{\top}g(F_{k}FS)$.
@@ B.3 Adaptive prediction sets
B.4 Motivation for test-time augmentation.
In practice, our trained orienter exhibits discontinuities at certain rotations. To demonstrate this, we perform an experiment on a bench shape, which is symmetric under a 180 degree rotation about the y-axis. We rotate the bench around this axis in increments of 1 degree and pass the resulting shape into our orienter for each rotation. We track two metrics throughout this process:
  • The quotient loss of our orienter’s prediction (i.e. the loss function from Problem 3, where we quotient the $L_{2}$ loss by the octahedral group). This measures the accuracy of our orienter’s predictions up to an octahedral symmetry.
  • The chamfer distance between the shapes obtained by applying the orienter’s output to the rotated bench at subsequent angles of rotation. This detects discontinuities in the orienter’s output.
We plot both of these metrics across the rotation angle in Figure 9. Our orienter exhibits several sharp discontinuities, manifested as spikes in the chamfer distance plot. These discontinuities are associated with spikes in the quotient loss, which are occasionally large. However, these spikes in the quotient loss are highly localized, and our orienter performs well for the vast majority of rotations. These localized spikes motivate our use of TTA to improve our pipeline’s performance. We provide implementation details for TTA in Appendices B.1 and B.2.
Figure 9: Our orienter exhibits several sharp discontinuities when rotating an input shape about a fixed axis (red curve). These discontinuities are associated with localized spikes in the quotient loss (blue curve), which motivate our use of TTA to improve our pipeline’s performance.